# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2

### v2

@@ 4.2 Sensor Data Understanding
We train SensorLLM on all five datasets using the same following parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for the other datasets.
@@
We assess SensorLLM’s ability to generate trend descriptions from sensor data, comparing it with the advanced GPT-4o 111gpt-4o-2024-08-06(OpenAI, 2024) to evaluate Sensor-Language Alignment. GPT-4o generates responses based on a predefined prompt (Appendix A.4), following our text template. We employ three evaluation methods:
  • NLP Metrics. We measure surface-level similarity and n-gram overlap using BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005). We also use SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021) to evaluate semantic similarity between model outputs and ground truth.
@@ Results.
Table 1 compares SensorLLM and GPT-4o on the Sensor Data Trend Analysis tasks, showing that our model consistently outperforms GPT-4o across all metrics. BLEU-1, ROUGE-1, ROUGE-L, and METEOR primarily focus on surface-level lexical or n-gram overlaps and SBERT and SimCSE can capture factual correctness or deeper semantic similarities. Across all metrics, SensorLLM generates trend descriptions more closely aligned with the ground truth. GPT-4o evaluations further highlight SensorLLM’s superior ability to capture trend details and coherence, whereas GPT-4o struggles with complex numerical data and trend observations (Yehudai et al., 2024). Human evaluation also favors SensorLLM, particularly for shorter sequences. CAPTURE-24 results are weaker compared to other datasets, likely due to its longer sequences being trained with the same parameters. Overall, these findings validate the effectiveness of our Sensor-Language Alignment method in enhancing LLMs’ ability to interpret complex numerical sequence. Appendix A.9 provides qualitative examples of outputs from both models.
Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
PatchTST 45.2±$\pm 1.48$ 86.8±$\pm 0.84$ 82.0±$\pm 0.71$ 80.0±$\pm 1.58$ 35.6±$\pm 0.89$
Ns-Transformer 52.6±$\pm 2.30$ 88.0±$\pm 0.71$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 34.8±$\pm 1.10$
Informer 51.2±$\pm 1.30$ 86.6±$\pm 1.14$ 78.0±$\pm 1.58$ 74.0±$\pm 0.71$ 35.6±$\pm 0.55$
Transformer 49.6±$\pm 1.67$ 85.4±$\pm 0.89$ 77.0±$\pm 0.71$ 75.2±$\pm 1.30$ 32.8±$\pm 0.84$
iTransformer 48.4±$\pm 1.82$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 80.4±$\pm 1.14$ 19.8±$\pm 0.84$
TimesNet 52.2±$\pm 2.39$ 87.4±$\pm 1.14$ 76.2±$\pm 1.92$ 78.4±$\pm 1.52$ 34.8±$\pm 0.84$
GPT4TS 54.2±$\pm 2.05$ 88.2±$\pm 0.84$ 80.4±$\pm 0.89$ 76.4±$\pm 1.14$ 32.8±$\pm 1.10$
Chronos+MLP 44.2±$\pm 1.30$ 82.2±$\pm 0.84$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 38.0±$\pm 0.71$
DeepConvLSTM 48.8±$\pm 2.39$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 75.0±$\pm 1.87$ 40.4±$\pm 0.89$
DeepConvLSTMAtt 54.0±$\pm 2.12$ 89.6 ±$\pm 1.14$ 79.2±$\pm 1.30$ 77.4±$\pm 2.19$ 41.4±$\pm 0.55$
Attend 60.2±$\pm 2.17$ 93.2±$\pm 0.84$ 84.6±$\pm 1.14$ 83.4±$\pm 1.14$ 43.6±$\pm 0.55$
SensorLLM 61.2±$\pm 3.56$ 91.2±$\pm 1.48$ 86.2±$\pm 1.48$ 89.4±$\pm 3.85$ 48.6±$\pm 1.14$
Table 2: F1-macro results (%) for the Task-Aware Tuning Stage, presented as the mean and standard deviation over 5 random repetitions. The top results for each dataset are highlighted as follows: Bold for the best and underline for the second-best.
@@ 4.3 Human Activity Recognition
In this section, we evaluate the performance of SensorLLM on HAR tasks. Each experiment runs for five trials, using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096. We report the F1 macro score A.8 to account for class imbalance across different activity categories.
@@ Results.
Table 2 reports the macro F1 scores (%) averaged over five random runs. SensorLLM achieves the highest performance on four datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24) and ranks second on UCI-HAR, demonstrating its effectiveness in handling diverse sensor data.
Notably, SensorLLM achieves a notable improvement on the CAPTURE-24 dataset, surpassing all baselines by a significant margin with a mean F1-macro score of 48.6%, which is 5.0% higher than Attend (43.6%). On USC-HAD, SensorLLM achieves the highest score of 61.2%, outperforming Attend, the second-best baseline, by 1.0%. Similarly, on PAMAP2, SensorLLM achieves a score of 86.2%, exceeding Attend (84.6%) by 1.6%. On MHealth, SensorLLM sets a new state-of-the-art with a score of 89.4%, surpassing Attend (83.4%) by 6.0%. These results highlight SensorLLM’s ability to consistently outperform existing methods across diverse datasets.
For UCI-HAR, SensorLLM achieves the second-best score (91.2%), slightly trailing Attend (93.2%). In contrast, Chronos+MLP shows only a slight improvement over iTransformer, the lowest-performing baseline (82.2% vs. 81.8%), indicating that Chronos embeddings alone have limited utility for HAR on this dataset. However, our framework significantly enhances their effectiveness, highlighting the robustness of our alignment approach.
@@ 5 Ablation Studies
USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\mathbf{\pm 3.56}$
UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\mathbf{\pm 1.48}$
PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\mathbf{\pm 1.48}$
MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\mathbf{\pm 3.85}$
CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\mathbf{\pm 1.14}$
Table 3: The results for SensorLLM trained with/without text prompts. Task-only refers to conducting HAR directly bypassing sensor-language alignment.
Impact of Alignment.
To evaluate the impact of alignment, we included the Chronos+MLP baseline in Section 4.3 to show that SensorLLM’s performance is not solely driven by Chronos embeddings. Additionally, we compared SensorLLM with a Task-only model, which skips the Sensor-Language Alignment Stage and directly applies Chronos embeddings and the LLM for HAR. As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of additional textual input. Notably, the Task-only model often performs on par with or worse than traditional TS baselines, underscoring the importance of our alignment method. These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance and that alignment is essential for enabling LLMs to effectively understand sensor data.
Impact of Prompts.
To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM’s performance with and without prompts. As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture. This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns. The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy. More broadly, the ability to process both sensor signals and textual prompts not only enhances classification performance but also extends LLMs’ potential for tackling complex sensor-driven tasks in future applications.
@@
Alignment Module Layers.
We investigate the effect of hidden layer depth in the alignment module MLP on UCI-HAR, PAMAP2, and MHealth. As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) led to mixed results. The average F1-macro scores improved on UCI-HAR (91.2% → 92.0%) and MHealth (89.4% → 90.2%), but slightly dropped on PAMAP2 (86.2% → 85.8%). These results suggest that the optimal number of hidden layers varies by dataset.
Effect of Model Size on Performance.
To assess the impact of model size, we tested SensorLLM-3b, a resource-efficient variant using Chronos-base and Llama3.2-3b. Experiments were conducted on USC-HAD, UCI-HAR, and MHealth. As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b across all datasets, illustrating the trade-off between model size and accuracy. However, SensorLLM-3b remains competitive, outperforming Attend on USC-HAD and MHealth, while trailing it only on UCI-HAR. These results indicate that SensorLLM-3b offers a resource-efficient alternative, maintaining strong performance relative to other baselines.
@@
Cross-Dataset Generalization.
To assess the robustness of SensorLLM, we conducted cross-dataset experiments, training the Sensor-Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa. These datasets share the same sensor channels but differ in sampling rates. As shown in Table 4, SensorLLM achieves performance comparable to training both stages on the same dataset. This indicates that once modality alignment is established, retraining on new datasets is unnecessary. SensorLLM effectively generalizes to downstream tasks across diverse datasets, demonstrating that the alignment stage enables the LLM to truly understand sensor data, rather than memorizing dataset-specific features. These results highlight SensorLLM’s potential for broad cross-dataset generalization, laying the groundwork for future TS-LLM models.
Stage 1 Stage 2 Results
USC-HAD UCI-HAR 91.0±$\mathbf{\pm 1.41}$
UCI-HAR USC-HAD 61.6±$\mathbf{\pm 2.07}$
Table 4: Cross-dataset experiments.
@@ 6 Conclusions
We introduced SensorLLM, a multimodal framework that aligns sensor data with automatically generated text at a human-perception level, moving beyond machine-level alignment. SensorLLM effectively captures complex sensor patterns, achieving superior performance in HAR tasks. Experiments across diverse datasets demonstrate its robustness in handling variable-length sequences, multi-channel inputs, and textual metadata. Cross-dataset results further highlight its strong generalizability without requiring dataset-specific alignment. This work establishes a foundation for Sensor-Text MLLMs, with potential applications for sensor data analysis. We release our code and data generation pipeline to facilitate future research on integrating time-series and text, particularly in low-resource domains.
@@ 7 Limitations
To ensure fair comparisons with existing HAR models (classification models), we adopt a classifier for downstream tasks rather than fully exploiting the LLM’s generative abilities. Although our results verify that the Sensor-Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may constrain the model’s adaptability to new activity categories. Future work could explore using the LLM in a generative or prompt-based capacity, enabling broader application scenarios such as activity discovery or open-set recognition.
@@ A.1 More related work
A.2 Data Generation
We generate text data from sensor readings using predefined sentence templates (5, Tables 6,, 7). These templates are randomly selected to create diverse question-answer (QA) pairs. To enhance variability, we employ GPT-4o to generate synonymous variations. Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
@@ A.6 Datasets
A.7 Baselines for Task-Aware Tuning Stage
@@
A.8 Evaluation Metrics for Task-Aware Tuning Stage
@@
A.9 Sensor-Language Alignment Stage Output Examples

### v3

@@ 4.2 Sensor Data Understanding
All datasets are trained using the same parameters in the Sensor-Language Alignment Stage: a learning rate of 2e-3, 8 epochs, batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 8192 for CAPTURE-24 and 4096 for others.
@@
We assess the performance of SensorLLM in the sensor–language alignment stage by comparing its ability to generate trend descriptions from sensor data with that of the advanced GPT-4o 111gpt-4o-2024-08-06 (OpenAI, 2024). GPT-4o generates responses using a predefined prompt (Appendix A.4). We adopt three evaluation methods:
  • NLP Metrics. We use BLEU-1 (Papineni et al., 2002), ROUGE-1, ROUGE-L (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to measure surface-level similarity and n-gram overlap. For deeper semantic alignment and factual correctness, we adopt SBERT (Reimers and Gurevych, 2019) and SimCSE (Gao et al., 2021).
@@ Results.
Table 1 compares SensorLLM and GPT-4o on the Sensor Data Understanding task. SensorLLM consistently outperforms GPT-4o across all metrics, with outputs that align more closely with ground truth and exhibit stronger trend understanding and coherence. In contrast, GPT-4o often struggles with complex numerical data and trend detection (Yehudai et al., 2024). When serving as an evaluator, GPT-4o also indicates a preference for SensorLLM. Human evaluation, conducted on shorter sequences, likewise favors SensorLLM. Performance on CAPTURE-24 is comparatively lower, likely due to longer sequences being trained under fixed parameters. Overall, these results validate the effectiveness of our alignment method in enabling LLMs to interpret complex TS data. Qualitative examples are provided in Appendix A.10.
@@ 4.3 Human Activity Recognition
In this section, we evaluate the performance of SensorLLM on HAR tasks. Each experiment is run for five trials using 8 training epochs, a batch size of 4, gradient accumulation steps of 8, and a maximum sequence length of 4096. We report both F1-macro (Appendix A.9) and accuracy to account for class imbalance and overall prediction performance across different activity categories.
@@ Results.
Table 2 reports F1-macro and accuracy scores (%) averaged over five runs. SensorLLM achieves the best performance on four out of five datasets (USC-HAD, PAMAP2, MHealth, CAPTURE-24), and ranks second on UCI-HAR, slightly behind Attend. It shows notable gains on challenging datasets such as CAPTURE-24 and MHealth, demonstrating strong performance in real-world and long-sequence settings. Compared to Chronos+MLP, which uses the same TS encoder, SensorLLM significantly improves both F1-macro and accuracy, highlighting the effectiveness of our alignment strategy in enabling LLMs to understand and classify sensor data.
Strong results on both F1-macro and accuracy indicate that SensorLLM performs well in both overall prediction and per-class balance, showing robust generalization across diverse sensor configurations, activity types, and data collection environments.
@@ 5 Ablation Studies
USC-HAD 43.4±$\pm 2.88$ 45.0±$\pm 1.58$ 49.6±$\pm 1.67$ 61.2±$\pm 3.56$
UCI-HAR 80.0±$\pm 2.12$ 82.0±$\pm 1.58$ 89.2±$\pm 1.10$ 91.2±$\pm 1.48$
PAMAP2 74.2±$\pm 2.28$ 75.4±$\pm 3.05$ 83.0±$\pm 0.71$ 86.2±$\pm 1.48$
MHealth 76.6±$\pm 1.34$ 77.4±$\pm 3.13$ 86.6±$\pm 1.14$ 89.4±$\pm 3.85$
CAPTURE-24 44.8±$\pm 0.84$ 46.0±$\pm 0.71$ 47.2±$\pm 0.84$ 48.6±$\pm 1.14$
Table 3: F1-macro scores for models trained with and without text prompts. Task-only refers to conducting Task-Aware Tuning directly bypassing the alignment stage.
Removing Alignment Hurts.
To assess the role of sensor–language alignment, we include the Chronos+MLP baseline (Section 4.3) to demonstrate that SensorLLM ’s performance is not solely due to the strength of the Chronos encoder. We further compare SensorLLM with a Task-only variant that skips the Sensor–Language Alignment stage and directly feeds Chronos embeddings into the LLM for HAR. As shown in Table 3, SensorLLM consistently outperforms the Task-only model across all five datasets, regardless of whether textual prompts are included. Notably, the Task-only model often performs comparably to or worse than traditional TS baselines, underscoring the critical role of alignment. These results confirm that Chronos embeddings alone are insufficient for optimal HAR performance, and that our alignment stage is essential for enabling the LLM to effectively interpret sensor data.
Textual Prompts Enhance HAR.
To assess the role of additional textual information (e.g., statistical features for each sensor channel) in the Task-Aware Tuning Stage, we compared SensorLLM ’s performance with and without prompts. As shown in Table 3, incorporating prompts consistently improves F1-macro scores across all datasets, with a more pronounced effect in the full SensorLLM architecture. This demonstrates that the model effectively integrates sensor and textual data, enhancing its ability to capture complex temporal patterns. The results highlight the benefits of multimodal inputs, which enrich sensor data representations and improve HAR accuracy. More broadly, the ability to jointly process sensor data and textual prompts underscores the potential of LLMs for more generalizable and interpretable sensor-driven applications.
@@
MLP Depth Trade-offs.
We examine how the depth of the alignment module MLP affects performance on UCI-HAR, PAMAP2, and MHealth. As shown in Figure 3, increasing the number of hidden layers from one ($1024\to 2048\to 4096$) to two ($1024\to 2048\to 3072\to 4096$) yields mixed results. F1-macro scores improve on UCI-HAR and MHealth, but slightly decrease on PAMAP2. These findings suggest that deeper MLPs do not always improve performance, and a single hidden layer offers a good balance between accuracy and efficiency.
Smaller SensorLLM Still Compete.
To address computational feasibility for deployment in resource-constrained environments, we evaluate SensorLLM-3b—a lighter variant built with Chronos-base and LLaMA3.2-3b. Experiments were conducted on USC-HAD, UCI-HAR, and MHealth. As shown in Figure 4, SensorLLM-3b achieves slightly lower performance than SensorLLM-8b, reflecting the trade-off between model size and accuracy. Nevertheless, it remains competitive—outperforming Attend on USC-HAD and MHealth, and closely trailing it on UCI-HAR. These results suggest that SensorLLM-3b provides a strong balance between efficiency and performance, making it a viable choice for real-world, resource-limited applications.
@@
Dataset F1-macro # Channels
w/o ST w/ ST
MHealth 89.6±$\pm 2.70$ 90.2±$\pm 3.11$ 15
PAMAP2 84.4±$\pm 1.14$ 85.8±$\pm 0.84$ 27
Table 4: Effect of special tokens on HAR based on two-layer alignment MLP. ST refers to special tokens.
Special Tokens Improve Performance.
We investigate the role of special tokens in helping SensorLLM distinguish sensor data from text and identify different sensor channel types. Special tokens are added to the aligned embeddings of each sensor channel and act as learned identifiers. They provide structural cues that help the LLM model channel-wise dependencies and reduce modality confusion. We conduct experiments on PAMAP2 and MHealth, both of which contain multiple sensor channels. As shown in Table 4, removing special tokens leads to a slight drop in F1-macro scores, with the performance gap tending to widen as the number of sensor channels increases. This confirms their value in preserving positional and channel-level structure within a flat token sequence.
Alignment Enables Generalization.
To assess the robustness of SensorLLM, we conduct cross-dataset experiments by training the Sensor–Language Alignment Stage on USC-HAD and the Task-Aware Tuning Stage on UCI-HAR, and vice versa. While these datasets share the same sensor channels, they differ in sensor wearing position, sampling rates and activity distributions. As shown in Table 5, SensorLLM achieves performance comparable to models trained entirely on the same dataset. This suggests that once modality alignment is learned, it can be transferred across datasets without retraining. These results indicate that SensorLLM does not overfit to dataset-specific patterns but learns generalizable sensor-language representations, demonstrating strong cross-dataset adaptability and paving the way for more universal TS–LLM frameworks.
Stage 2 Stage 1 F1-macro
USC-HAD UCI-HAR 61.6±$\pm 2.07$
USC-HAD 61.2±$\pm 3.56$
UCI-HAR UCI-HAR 91.2±$\pm 1.48$
USC-HAD 91.0±$\pm 1.41$
Table 5: F1-macro scores for cross-dataset experiments.
@@ 6 Conclusions
We present SensorLLM, a multimodal framework that aligns sensor data with natural language through a QA format at a human-perception level, moving beyond machine-level alignment. It effectively captures complex sensor patterns, achieves strong performance on HAR tasks, and generalizes well without requiring dataset-specific alignment. Experiments demonstrate its robustness across variable-length sequences, multivariate inputs, and textual metadata. To support future research, we release our code and the SensorQA dataset, constructed from five public HAR benchmarks, to advance time-series and language integration, particularly in low-resource domains.
@@ 7 Limitations
To ensure fair comparisons with existing HAR models, we adopt a classifier for downstream tasks rather than fully leveraging the LLM’s generative capabilities. While our results demonstrate that the Sensor–Language Alignment Stage can generalize across datasets, relying on a fixed-class classifier may limit adaptability to new activity categories. Although zero-shot adaptability is a valuable direction, we did not explore it here due to the lack of comparable baselines. To the best of our knowledge, no prior work supports generalization to unseen activity classes under variable-length and variable-channel sensor input, as our framework does, making fair comparison on zero-shot settings infeasible at this stage. Future work could explore generative or prompt-based approaches to support broader applications such as activity discovery or open-set recognition.
@@ A.1 More related work
A.2 SensorQA Generation
For SensorQA, we generate text data from sensor readings using predefined sentence templates (Tables 6, 7, 8). These templates are randomly selected to create diverse question-answer (QA) pairs. To enhance variability, we employ GPT-4o to generate synonymous variations. Each sentence contains placeholders for numerical values (e.g., timestamps, sensor readings) or textual information, which are dynamically replaced to produce coherent QA pairs aligned with the sensor data.
@@ A.6 Datasets
A.7 SensorQA
We introduce SensorQA, a novel question-answering dataset designed to align time-series sensor data with human-interpretable natural language. Each sample pairs a segment of sensor input with natural language questions and answers, capturing trends, patterns, or activity-level semantics.
Sensor-Language Alignment Stage
of SensorQA focuses on aligning uni-variate sensor sequence of variable length with descriptive textual responses and includes two types of QA tasks:
  • Trend Analysis QA, which describes how the signal changes within the window.
  • Trend Summary QA, which summarizes the overall behavior across a window in a concise natural language phrase.
Task-Aware Tuning Stage
focuses on using multi-variate sensor sequences to perform human activity classification, leveraging the aligned representations learned in the alignment stage. This stage of SensorQA contains statistical information from each sensor channel as part of the input representation.
Dataset Stage 1 Stage 2
Train Test Train Test
USC-HAD 300,744 58,704 22,790 4,555
UCI-HAR 128,292 25,932 7,352 2,947
PAMAP2 738,666 271,674 14,163 5,210
MHealth 283,020 60,780 4771 2,039
CAPTURE-24 72,714 35,688 61,327 30,138
Table 12: Training and testing sample counts for Stage 1 and Stage 2 across datasets of SensorQA.
The distribution of training and testing data across both stages is summarized in Table 12.
A.8 Baselines for Task-Aware Tuning Stage
@@
A.9 Evaluation Metrics for Task-Aware Tuning Stage
@@
A.10 Sensor-Language Alignment Stage Output Examples